# hlinnerleftmargin=0, innerrightmargin=0, innertopmargin=0, innerbottommargin=0, backgroundcolor=lightyellow, linewidth=0 Ultra-Efficient Thermally Undercut Silicon Photonic Devices in a 300 mm CMOS Foundry Process

### v1

@@ Discussion
In conclusion, we demonstrate a wafer-scale thermal undercut in a 300 mm CMOS foundry applied to a range of fundamental silicon photonic devices with state-of-the-art performance and yield. We show an excellent 40 improvement in tuning efficiency of linear thermo-optic phase shifters with undercut, enabling a drastic reduction in power consumption of large-scale silicon photonic circuits. Undercut microring and racetrack-style modulators also exhibit high tuning efficiency, owing to their ability to accommodate both larger and elongated external trench openings along with an internal trench opening. Further, we apply the developed undercut etch to state-of-the-art microdisk modulators, achieving an almost 5 improvement in tuning efficiency. These efficiently tunable compact vertical-junction microdisk modulators with high modulation efficiency are fully compatible with CMOS driving voltages, enabling scalable, ultra efficient DWDM links.
By systematically varying the undercut trench geometry and measuring tuning efficiencies across multiple full wafers of microdisk modulators, we identified key trends that reveal how trench dimensions impact thermal tuning efficiency. Wider and more elongated trenches consistently provide improved thermal isolation, resulting in significant reductions in the power required for thermal tuning. Further, reduced trench separation also resulted in higher tuning efficiency. This finding suggests that trench geometry optimization can serve as a powerful design parameter for further minimizing power consumption in future device designs. Notably, through full wafer-scale measurements across 300 mm wafers, we additionally demonstrate the robustness of the thermally undercut devices through the high yield and low standard deviation. By establishing predictable efficiency gains based on trench geometry, we provide a robust framework for integrating thermal undercut as a standardized process across different CMOS fabrication lines, ensuring reproducibility and consistency at the scale required for high-volume production. This level of scalability is critical for enabling low-cost, high-throughput manufacturing of photonic integrated circuits where consistent thermal performance is essential.
@@ Acknowledgements
This work was supported in part by the U.S. Advanced Research Projects Agency–Energy under ENLITENED Grant DE-AR000843, in part by the U.S. Defense Advanced Research Projects Agency under PIPES Grant HR00111920014, and in part by the Center for Ubiquitous Connectivity (CUbiC), sponsored by the Semiconductor Research Corporation (SRC) and DARPA under the JUMP 2.0 program. This material is based on research sponsored by the United State Air Force (FA8750-23-C-1001) and Air Force Research Laboratory under AIM Photonics (agreement number FA8650-21-2-1000) and also FA8750-21-2-0004. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of the United States Air Force, the Air Force Research Laboratory or the U.S. Government. Approved for Public Release; Distribution A Unlimited: AFRL-2025-1217.

### v2

@@ Discussion
In conclusion, we developed and experimentally verified a wafer-scale thermal undercut process in a 300 mm CMOS foundry applied to a range of fundamental silicon photonic devices with state-of-the-art performance and yield. We show an excellent 40 improvement in tuning efficiency of linear thermo-optic phase shifters with undercut, enabling a drastic reduction in power consumption of large-scale silicon photonic circuits. Undercut microring and racetrack-style modulators also exhibit high tuning efficiency, owing to their ability to accommodate both larger and elongated external trench openings along with an internal trench opening. Further, we apply the developed undercut etch to state-of-the-art microdisk modulators, achieving an almost 5 improvement in tuning efficiency. These efficiently tunable compact vertical-junction microdisk modulators with high modulation efficiency are fully compatible with CMOS driving voltages, enabling scalable, ultra efficient DWDM links. Through comprehensive wafer-scale measurements comparing nominal non-undercut and undercut devices, we show high uniformity in across entire 300 mm wafers, indicating high yield and reproducibility for the developed process. Notably, we observe that high nominal power consumption (large ) can be largely offset by thermally undercutting the device, greatly relaxing the heater design requirements and allowing the heater to be placed far from the optical mode.
Furthermore, by systematically varying the undercut trench geometry and measuring tuning efficiencies across multiple full wafers of microdisk modulators, we identified key trends that reveal how trench dimensions impact thermal tuning efficiency. Wider and more elongated trenches consistently provide improved thermal isolation, resulting in significant reductions in the power required for thermal tuning. Further, reduced trench separation also resulted in higher tuning efficiency. This finding suggests that trench geometry optimization can serve as a powerful design parameter for further minimizing power consumption in future device designs. By establishing predictable efficiency gains based on trench geometry, we provide a robust framework for integrating thermal undercut as a standardized process across different CMOS fabrication lines, ensuring reproducibility and consistency at the scale required for high-volume production. This level of scalability is critical for enabling low-cost, high-throughput manufacturing of photonic integrated circuits where consistent thermal performance is essential.
@@ Acknowledgements
This work was supported in part by the U.S. Advanced Research Projects Agency–Energy under ENLITENED Grant DE-AR000843, in part by the U.S. Defense Advanced Research Projects Agency under PIPES Grant HR00111920014, and in part by the Center for Ubiquitous Connectivity (CUbiC), sponsored by the Semiconductor Research Corporation (SRC) and DARPA under the JUMP 2.0 program. This material is based on research sponsored by the United States Air Force (FA8750-23-C-1001) and Air Force Research Laboratory under AIM Photonics (agreement number FA8650-21-2-1000) and also FA8750-21-2-0004. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of the United States Air Force, the Air Force Research Laboratory or the U.S. Government. Approved for Public Release; Distribution A Unlimited: AFRL-2025-1217.